A flood control project with a life of 21 years will require an investment of $150,000 and annual maintenance costs of $10,000. The project will provide no benefits for the first two years but will save $100,000 per year in flood damage starting in the third year. The appropriate MARR is 15% per year. What is the conventional B-C ratio with present worth for the flood control project. (Enter your answer rounded to two decimal places; i.e., x.xx)
Answer #1

Solution :- Present value of annual benefits = 100000 * Cumulative present value factor from year 3 to year 21 at 15 % (using present value table)

= 100000 * 4.687 (approx).

= $ 468700.

Present value of annual maintenance costs = 10000 * Cumulative present value factor for 21 years at 15 % (using present value table)

= 10000 * 6.312 (approx)

= $ 63120 (approx)

Present value of cash inflows = Present value of annual benefits - Present value of annual maintenance costs.

= 468700 - 63120

= $ 405580.

Present value of initial cash outflow = $ 150000 (Given in the question)

Conventional B-C ratio =  Present value of cash inflows /  Present value of initial cash outflow.

= 405580 / 150000.

= 2.70 (approx).

Conclusion :- Conventional B-C ratio = 2.70 (can be mentioned as 2.70 : 1)